the sum of 3 consecutive even numbers is equal to one less than the first consecutive even number, doubled.

Answer by checkley79(3341)

X+(X+2)+(X+4)=2(X/2-1)
3X+6=X-2
3X-X=-2-6
2X=-8
X=-8/2
X=-4 ANS. FIR THE FIRST EVEN NUMBER.
-4+2=-2 FOR THE SECOND EVEN NUMBER.
-4+4=0 FOR THE THIRD EVEN NUMBER.
PROOF:
-4+(-2)+0=2(-4/2-1)
-4-2=2(-2-1)
-6=2*-3
-6=-6